Joe and Vickie’s manager wants to forecast the number of sales that will require fulfillment in...

  1. Joe and Vickie’s manager wants to forecast the number of sales that will require fulfillment in the next five months. Below are the current pipeline activities:

Interest Creation

Pre-purchase

Purchase

Post-purchase

Salesperson

Cold Leads

Warm Leads

Prospects

1st/2nd Meeting

2nd/3rd Meeting

Delivery

Support

Joe

21

17

11

8

4

11

17

Vickie

38

19

16

9

6

3

19

These historical averages exist for each of the activities:

  • 3% of cold calls are converted to sales within five months.

  • 19% of warm calls are converted to sales within four months.

  • 31% of prospects are converted to sales within three months.

  • 47% of customers who agree to a pre-purchase meeting are converted to sales within two months.

  • 64% of customers who agree to a purchase meeting are converted to sales within one month.

What is the sales pipeline volume?
Answer #1
  • 3% of cold calls are converted to sales within five months = 3%*(21+38) = 1.77
  • 19% of warm calls are converted to sales within four months = 19%*(17+19) = 6.84
  • 31% of prospects are converted to sales within three months = 31%*(11+16) = 8.37
  • 47% of customers who agree to a pre-purchase meeting are converted to sales within two months = 47%*(8+9) = 7.99
  • 64% of customers who agree to a purchase meeting are converted to sales within one month = 64%*(4+6) = 6.4

Hence the total sales pipeline volume = 1.77 + 6.84 + 8.37 + 7.99 + 6.40 = 31.37

Hence sales pipeline volume is 31(rounding-off)